12 A nickel-titanium alloy is used to make components for jet turbine aircraft engines. Cracking is a potentially serious proHi, can you help me provide the step-by-step working solution for part (b) of the above question?
Answer #1

A titanium-nickel alloy is a superalloy. It is mostly used for making jet turbine aircraft engines. Jet turbines work from high cyclic dynamic, thermal loads are acting on the jet engine body. Therefore,  cracking is occurs in their parts and it is a very sober problem So, that we make an experimental test for four factors on cracks.

The four fours are pouring temperature (A), titanium (B), heat treatment (C), amount of grain refiner used (D). The best optimum technique can be a 24 Experiment design technique, which is developed by Dr. Taguchi.

The data are shown below in the design table and total of the replica I and II and the effect of factors

A B C D Replicate I Replicate II Total
-1 -1 -1 -1 7.037 6.376 13.413
1 -1 -1 -1 14.7.7 15.219 15.219
-1 1 -1 -1 11.635 12.089 23.724
1 1 -1 -1 17.273 17.815 35.088
-1 -1 1 -1 10.403 10.151 20.554
1 -1 1 -1 4.368 4.098 8.466
-1 1 1 -1 9.36 9.253 18.613
1 1 1 -1 13.44 12.923 26.363
-1 -1 -1 1 8.561 8.951 17.512
1 -1 -1 1 16.867 17.052 33.919
-1 1 -1 1 13.876 13.658 27.534
1 1 -1 1 19.824 19.639 39.463
-1 -1 1 1 11.846 12.337 24.183
1 -1 1 1 6.125 5.904 12.029
-1 1 1 1 11.19 10.935 22.125
1 1 1 1 15.653 15.053 30.706

A B AB CAC BC ABC D AD BD ABD CD ACD BCD ABCD TOTAL -1 -1 1 1 1 1 -1 -1 1 1 -1 1 -1 - 1 1 13.413 1 -1 -1 -1 -1 1 1 -1 -1 1 1

Effectr = Contrast/2k – 1)n = Contrast/8 * 2)

EffectA = 48.302/16 = 3.018875SSa = (48.302*48.302) 32 = 72.908850126Effect AB = 30.946/16 = 1.934126 SSAB = 30.946 * 30.946)/32 = 29.926716125

EffectABC = 50.2/16 = 3.1375SSAB = (50.2 * 50.2)/32 = 78.75125

Effect ABCD = 0.226/16 = 0.014125SS ABCD = (0.226 -0.226)/32 = 0.001596126

Interaction Plot for Replicate I & II Data Means

The Factors effect plot

Main Effects Plot for Replicate I & II Data Means А в 22.5-1 20.0 Mean 25.0 22.5 20.0

Fitted Line plot for Factor A

Fitted Line Plot A = -0.6622 +0.02872 Replicate I & II 1.0- . . 1.03631 6.0% R-59 R-Sq(adj) 0.0% -1.0 - 10 15 20 25 Replicate

Fitted Line plot for Factor B

Fitted Line Plot B = - 1.544 + 0.06696 Replicate I & II R-59 R-Sq(adj) 0.876503 32.8% 28.0% -1.0 - 10 .... 15 . . 20 25 RepliFitted Line plot for Factor C

Fitted Line Plot C = 0.8443 -0.03662 Replicate I & II .. 1.0 . . R-59 R-Sq(adj) 1.01529 9.8% 3.4% с -1.04 10 15 20 25 Replica

Fitted Line plot for Factor D

Fitted Line Plot D = -0.9074 + 0.03935 Replicate I & II 1.04 . . R-54 R-Sa(adj) 1.00671 11.3% 5.0% D 0.0- -1.0 . 10 .. 15 ..

Minitab Output

Fitted Line: D versus Replicate I & II


Regression Analysis: Replicate I & II versus A, B, C, D

The regression equation is
Replicate I & II = 23.1 + 2.10 A + 4.90 B - 2.68 C + 2.88 D


Predictor Coef SE Coef T P
Constant 23.057 1.632 14.13 0.000
A 2.100 1.632 1.29 0.225
B 4.895 1.632 3.00 0.012
C -2.677 1.632 -1.64 0.129
D 2.877 1.632 1.76 0.106


S = 6.52716 R-Sq = 59.9% R-Sq(adj) = 45.4%


Analysis of Variance

Source DF SS MS F P
Regression 4 701.02 175.26 4.11 0.028
Residual Error 11 468.64 42.60
Total 15 1169.66


Source DF Seq SS
A 1 70.54
B 1 383.39
C 1 114.67
D 1 132.43